trying to figure out how to work out the answer to (x - 2)squared + x squared = (x + 2)squared i worked it down to x squared - 8x = 0
Answer by fractalier(6550)

Okay we have
(x-2)^2 + x^2 = (x+2)^2
x^2 - 4x + 4 + x^2 = x^2 + 4x + 4
and yes
x^2 - 8x = 0
x(x - 8) = 0
and x = 0 or x = 8